OpenStudy (anonymous):

Multiply and simplify: (4 + v8) (4- v8)

OpenStudy (muhammad_nauman_umair):

is it v^8?

jhonyy9 (jhonyy9):

16-64v^2

jhonyy9 (jhonyy9):

16(1-4v^2)

OpenStudy (muhammad_nauman_umair):

as, a^2-b^2=(a+b)(a-b)

OpenStudy (anonymous):

FOIL (4+v8) (4- v8) =16 - 4v8 + 4v8 - 8 = 8

OpenStudy (muhammad_nauman_umair):

then, =4^2-(v^8)^2 =16-v^16

OpenStudy (anonymous):

@vi3tbloodxx you're wrong... (4+v8) (4- v8) =16 - 4v8 + 4v8 - 8 = 8 it should be (4+v8) (4- v8) =16 - 4v8 + 4v8 - 64v^2 =4^2-(8v)^2 so from here we can see that there's no need to multiply and simplify... just use difference of binom formula : a^2-b^2 =(a+b)(a-b) \[\LARGE a^2-b^2=(a-b)(a+b)\]

OpenStudy (anonymous):

a^2−b^2=(a−b)(a+b)\ in this equation (4+v8) (4- v8) a = 4 and b = v8 4^2 - (v8)^2 gives 16 - 8 = 8

OpenStudy (anonymous):

@vi3tbloodxx Still don't get what you're trying to say...O_o where's "v" how come (v8)^2=8 ???

OpenStudy (anonymous):

the v stands for the square root of 8. When squaring a square root the result is the number inside of the square root.

OpenStudy (anonymous):

v cannot be a variable due to the fact that variables stands behind the number. Also from the poster previous post, she been using "v" as a square root notation.

OpenStudy (anonymous):

hold on... is \[\LARGE v8=\sqrt8\] or \[\LARGE v8=v8\] ?? "v" is a square root or just a letter?? !

OpenStudy (anonymous):

auff... not your fault , I misunderstood ! I apologize , you're right , the answer is 8 :(